OpenStudy (anonymous):

What is the mean absolute deviation for 9, 7, 6, 8, 7, and 5? 1 2 6 7

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@tkhunny hello :D

OpenStudy (tkhunny):

Have you considered calculating the mean and the subsequent absolute deviations?

OpenStudy (anonymous):

Well i came here for help to learn how.. D:

OpenStudy (tkhunny):

I do not believe you cannot calculate the mean. Go!

OpenStudy (anonymous):

mean is 7

OpenStudy (tkhunny):

Perfect. Now, how far away from the mean is each of the values? We want ABSOLUTE deviation, so don't worry about direction. Just how far away from 7 is it?

OpenStudy (anonymous):

2 1 2 3 4 are the numbers i got if im understanding correctly

OpenStudy (tkhunny):

I'm not sure how you managed those values. 9 - 7 = 2 ==> 2 7 - 7 = 0 ==> 0 6 - 7 = -1 ==> 1 8 - 7 = 1 ==> 1 7 - 7 = 0 ==> 0 5 - 7 = -2 ==> 2 You should get 2 0 1 1 0 2

OpenStudy (anonymous):

ok and what do i do with the numbers you gave me?

OpenStudy (tkhunny):

Are you sure they are correct? What is an "Absolute Deviation"? Think on it and BE SURE!

OpenStudy (anonymous):

i meant 2 0 1 1 0 2

OpenStudy (tkhunny):

So did I. Do you believe those values? If you don't believe and aren't sure where they came from, we can't move on.

OpenStudy (anonymous):

they are correct.

OpenStudy (anonymous):

i looked at the math its correct.

OpenStudy (tkhunny):

Calculate the mean of those values.

OpenStudy (anonymous):

1

OpenStudy (tkhunny):

Review: 1) Calculate the mean of the data. 2) Calculate the deviation from the mean for each datum. This is the "Deviation". 3) Use "Absolute Value" on each deviation to create the "Absolute Deviation". 4) Find he mean of the values from #3 to create the "Mean Absolute Deviation". You should see how the very name of the calculation suggests how to calculate it. Good work.

OpenStudy (anonymous):

Thank you sensei! xD

OpenStudy (tkhunny):

I hope you have accomplished your stated goal, "Well i came here for help to learn how.. D:". :-)